# A Simple Latent Diffusion Approach for Panoptic Segmentation and Mask Inpainting

### v1

@@ Abstract
Panoptic and instance segmentation networks are often trained with specialized object detection modules, complex loss functions, and ad-hoc post-processing steps to handle the permutation-invariance of the instance masks. This work builds upon Stable Diffusion and proposes a latent diffusion approach for panoptic segmentation, resulting in a simple architecture which omits these complexities. Our training process consists of two steps: (1) training a shallow autoencoder to project the segmentation masks to latent space; (2) training a diffusion model to allow image-conditioned sampling in latent space. The use of a generative model unlocks the exploration of mask completion or inpainting, which has applications in interactive segmentation. The experimental validation yields promising results for both panoptic segmentation and mask inpainting. While not setting a new state-of-the-art, our model’s simplicity, generality, and mask completion capability are desirable properties. The code and models will be made available. [^1] https://github.com/segments-ai/latent-diffusion-segmentation
@@ 1 Introduction
The image segmentation task [47, 51] has gained a lot of attention in the literature, encompassing three popular subfields: semantic, instance, and panoptic segmentation. Over the years, segmentation tools have proven their usefulness for a wide range of applications, such as autonomous driving [19], medical imaging [50], agriculture [17], and augmented reality [1, 26]. Current methods are built upon convolutional networks [28] and transformers [73, 21] to learn hierarchical image representations, and to simultaneously leverage large-scale datasets [85, 45]. Earlier segmentation approaches relied on specialized architectures [29, 7, 67, 75], such as region proposal networks [58] and dynamic convolutions [34]. More recent approaches [16, 8] advocate an end-to-end strategy but introduce complex loss functions, i.e., bipartite matching using object queries to separate instances. Some works have shown promising results without the necessity for labels [70, 71, 72, 18, 76, 77], but likewise require highly-specialized modules, such as region proposal networks or clustering. Differently, we seek to leverage generative models to bypass the aforementioned components.
This observation aligns with recent works [49, 12, 41, 10, 11] advocating for general computer vision models in and attempt to unify the field. These pioneering works limit the adoption of task-specific components and instead use a generative process, such as autoregressive modeling, rendering its design applicable to different tasks and/or modalities. In a similar vein, we take inspiration from recent text-to-image diffusion models [59, 57, 52, 20, 31, 32, 56] to tackle the segmentation task in a generative fashion. In addition to the diffusion model’s general architecture, we further motivate this decision as follows: (i) diffusion models are capable of generating images with high photorealism and diversity; (ii) they perform on par with autoregressive priors while being more computationally efficient [63, 57]; (iii) they learn high-quality spatial representations, advantageous for dense prediction tasks; (iv) they naturally exhibit image editing capabilities. We build further upon these properties by harnessing the generative power of Latent Diffusion Models (LDMs) [59] for image segmentation applications.
@@
Figure 1: We present a simple generative approach for the segmentation task that builds upon Stable Diffusion [59]. The key idea is to leverage the diffusion process to bypass complex detection and segmentation modules, and to unlock mask inpainting. The generative process is conditioned on RGB images to predict panoptic masks in an iterative fashion.
@@
To realize this objective, we introduce LDMSeg, a simple Latent Diffusion Model for Segmentation, visualized in figure 1. In summary, our contributions are threefold:
@@
  • Generative Framework: This paper proposes a fully generative approach based on Latent Diffusion Models (LDMs) for panoptic segmentation. Our approach builds upon Stable Diffusion [59] to strive for simplicity and to ease compute. We first study the class-agnostic setup to liberate panoptic segmentation from predefined classes [40].
@@
  • General-Purpose Design: Our approach circumvents specialized architectures, complex loss functions, and object detection modules, present in the majority of prevailing methods. Here, the denoising objective omits the necessity for object queries, region proposals, and Hungarian matching [42]. This simple and general approach paves the way for future extensions to a wide range of dense prediction tasks, e.g., depth prediction, saliency estimation, etc.
@@
  • Mask Inpainting: We successfully apply our approach to scene-centric datasets and demonstrate its mask inpainting capabilities for different sparsity levels.
@@ 2 Related Work
Panoptic segmentation [38] has lately gained popularity as it combines semantic and instance segmentation. In particular, its goal is to detect and segment both stuff-like (e.g., vegetation, sky, mountains, etc.) and thing-like (e.g., person, cat, car, etc.) categories. Earlier works modified instance segmentation architectures to additionally handle (amorphous) stuff categories as they are hard to capture with bounding boxes. For instance, Kirillov et al. showed promising results by making independent predictions using semantic and instance segmentation architectures [38], and later by integrating a semantic segmentation branch with a feature pyramid network (FPN) into Mask R-CNN [37]. Other works [79, 14] extend this idea by relying on specialized architectures and loss functions from both segmentation fields. More recent works [8, 74, 16, 15, 81, 84] handle things and stuff categories in a unified way via object queries and Hungarian matching [42]. These works generally depend on specialized modules – such as anchor boxes [58], non-maximum suppression [33, 6], merging heuristics [38, 79], or bipartite matching algorithms [8, 16, 74], etc. – to generate panoptic masks. Instead, we propose a task-agnostic generative framework to bypass these components. Consequently, we refrain from using task-specific augmentations, such as large-scale jittering or copy-paste augmentations [25].
General-purpose Frameworks.
@@
Similar to our work, a number of task-agnostic solutions have been suggested that cast vision problems as a generative process. Kolesnikov et al. [41] minimized task-specific knowledge by learning task-specific guiding codes, using a language model, to train a separate vision model. Chen et al. [11] simplified this procedure by framing several vision tasks as language modeling tasks, within a single model. Lu et al. [49] also follow this route and show promising results for a large variety of vision and language tasks using a unified framework. Each work presents the input as a sequence of discrete tokens which are subsequently reconstructed via autoregressive modeling. Other works [78, 4] leverage masked image modeling to train a single model for multiple vision tasks. Differently, we leverage the denoising process in continuous latent space, which is well-suited to handle dense prediction tasks with high-dimensional inputs [32, 12, 59].
Denoising Diffusion Models.
@@
Denoising diffusion models [62, 32, 63] were introduced as a new class of generative models. Recent strategies [20, 52, 31, 57] additionally leverage text as guidance, e.g., via CLIP embeddings, to achieve results with impressive realism and control. Building upon its success, a few diffusion-based solutions appeared in the segmentation literature. However, they have undesirable properties: (i) the inability to differentiate between instances in an image [2, 3, 5], (ii) the necessity for specialized architectures and loss functions [27, 80], or (iii) the dependence on object detection weights and bit diffusion [12, 13]. The closest related work from Chen et al. [12], presents a framework for panoptic segmentation by leveraging the diffusion process in pixel-space. In contrast, we rigorously follow latent diffusion models by relying on continuous latent codes without the necessity for object detection modules. Importantly, we can effortlessly leverage public image-diffusion weights [59] as we keep the architecture task-agnostic.
@@ 3 Method
Figure 2: Overview of LDMSeg. Inspired by latent diffusion models, we present a simple diffusion framework for segmentation and mask inpainting. The approach consists of two stages: (i) learn continuous codes $z_{t}$ with a shallow autoencoder on the labels (Sec. 3.1); (ii) learn a denoising function conditioned on image latents $z_{i}$ (Sec. 3.2). In the second stage, the error between the predicted noise $\hat{\epsilon}$ and the applied Gaussian noise $\epsilon$ is minimized. During inference, we traverse the denoising process by starting from Gaussian noise. The models $f_{t}$ and $f_{i}$ respectively encode the labels and images. While we rely on the image encoder $f_{i}$ from Stable Diffusion [59], we focus on $f_{t}$ and $g$ for segmentation. Our aim is to prioritize generality by limiting task-specific components.
Preliminaries.
The aim of this paper is to train a fully generative model for panoptic segmentation via the denoising diffusion paradigm [32]. While this generative approach results in longer sampling times than discriminative methods, we justify this decision through four different lenses: (i) ease of use – diffusion models omit specialized modules, such as non-maximum suppression or region proposal networks, are stable to train, and exhibit faster sampling than autoregressive models [68]; (ii) compositionality – this generative approach captures complex scene compositions with high realism and diversity, while also enabling image editing [57]; (iii) dataset-agnostic – we rely on spatially structured representations that are not tied to predefined classes or taxonomies; (iv) computational cost – latent diffusion models [59] reduce the computational requirements by modeling latents instead of pixels with an autoencoder [36, 69]. Motivated by these points, we create a framework that builds upon latent diffusion models’ generative power [59, 57] for the segmentation task.
Problem Setup.
The considered segmentation task requires a dataset of images $\mathcal{X}=\{x_{1},\ldots,x_{n}\}$ and corresponding ground truth panoptic segmentation masks $\mathcal{Y}=\{y_{1},\ldots,y_{n}\}$. We don’t make any distinction between things or stuff classes but handle all classes identically. Particularly, we first only predict the individual panoptic IDs and not the individual classes, rendering our approach class-agnostic. We will also show results in the experiments when predicting class-labels.
@@
From a high-level perspective, our method has two key components: learning the prior over segmentation latents $z_{t}$ and learning a conditional diffusion process in latent space. First, we train a shallow autoencoder to capture the prior distribution $p(z_{t})$ that learns to compress the labels into compact latent codes $z_{t}$. Second, we train a diffusion process – conditioned on the image and target features $p(y|z_{t},z_{i})$. This component is responsible for decoding noisy target features, guided by image features $z_{i}$. We make a similar derivation as [57] but condition the generative process on images. Formally, this two-step procedure allows us to construct the conditional distribution $p(y|x)$ via the chain rule:
$p(y|x)=p(y|z_{t},z_{i})\cdot p(z_{t}).$ (1)
Both terms $p(z_{t})$ and $p(y|z_{t},z_{i})$ are reflected in our framework as two separate training stages (see figure 2). Section 3.1 discusses the learning of the prior via autoencoding and section 3.2 leverages this prior to train a latent diffusion model for segmentation.
3.1 Stage 1: Compress Segmentation Targets
In the first step, we train a network to compress the task-specific targets into latent codes. While we will focus on panoptic segmentation masks, we note that a similar strategy is applicable to other dense prediction targets, such as depth or surface normals estimation.
Motivation.
The motivation to design a shallow autoencoder stems from the observation that segmentation maps differ fundamentally from images as they are lower in entropy. First, these masks typically contain only a small number of unique values as they only capture the object’s general shape and location in the scene. Second, neighbouring pixels are strongly correlated and often identical, resulting in largely spatially redundant information. We conclude that the segmentation task only necessitates a shallow network to efficiently compress the task’s targets and to reliably capture the prior distribution $p(z_{t})$. Interestingly, this conclusion holds for a myriad of dense prediction tasks, such as panoptic segmentation, depth prediction, saliency estimation etc. This also justifies why we refrain from using more advanced autoencoders that rely on computationally demanding architectures with adversarial or perceptual losses [35, 83], e.g., VQGAN [23], typically used to encode images [59].
Encoding.
We analyze several encoding strategies to represent the input segmentation map $y$, latents $z_{t}$ and output $\hat{y}$.
1. Input $y$: Let $N$ denote the maximum number of instances per image. RGB-encoding (3 channels), bit-encoding ($\log_{2}N$ channels), one-hot-encoding ($N$ channels), or positional-encoding [73] are all justifiable options. We follow [12] in representing the instance IDs as bits, which is a natural choice given discrete segmentation maps.
@@ 3.1 Stage 1: Compress Segmentation Targets
Architectural Design.
The architecture of the autoencoder follows a simple and shallow design. It comprises an encoder $f_{t}$ and decoder $g$, yielding the overall function $g\circ f_{t}$. The encoder $f_{t}$ includes only a few strided convolutions to compress the targets and is inspired by ControlNet [82]. For instance, targets with size 512$\times$512 can be resized efficiently to 64$\times$64, in order to leverage the latent space of Stable Diffusion [59] by stacking 3 convolutions of stride 2. Similarly, the decoder $g$ consists of one or more transpose convolution to upscale the masks and minimize a loss at pixel-level. As a consequence, the number of trainable parameters is at least 2 orders of magnitudes smaller than the amount of parameters in the diffusion model $h$ ($\approx 2$M vs. $800$M). The model’s shallow design brings several advantages to the table: fast training, good generalization across datasets, and applicable to inpainting without architectural changes.
Loss Function.
The autoencoder aims to minimize the reconstruction error $\mathcal{L}_{\text{rec}}$ between the outputs $\hat{y}$ and the one-hot encoded segmentation masks $y$. While regression losses are a valid choice, we opt for a categorical loss due to segmentation task’s discrete nature. The reconstruction loss comprises two loss terms: (i) the cross-entropy loss $\mathcal{L}_{\text{ce}}$ enforces unique and confident predictions for each pixel; (ii) the mask loss $\mathcal{L}_{\text{m}}$ further refines the segmentation masks by treating each instance individually, alleviating the need for exhaustively labeled images. This term is implemented via the BCE and Dice losses [66, 29]. Note that this autoencoding strategy prevents the need for Hungarian matching [42].
Typically, latent diffusion models incorporate a penalty term $\Omega$ into the loss formulation to align the bottleneck latents with a standard Gaussian distribution $\mathcal{N}(0,I)$. This term generally takes the form of a KL divergence, resulting in a variational autoencoder [36]. However, we empirically observed that weight decay regularization suffices to keep the weights $w$, and by extension the latents $z_{t}$, bounded. Consequently, the magnitude of the model’s weights $||w||_{2}$ is penalized, resulting in the final loss:

### v2

@@ Abstract
Panoptic and instance segmentation networks are often trained with specialized object detection modules, complex loss functions, and ad-hoc post-processing steps to manage the permutation-invariance of the instance masks. This work builds upon Stable Diffusion and proposes a latent diffusion approach for panoptic segmentation, resulting in a simple architecture that omits these complexities. Our training consists of two steps: (1) training a shallow autoencoder to project the segmentation masks to latent space; (2) training a diffusion model to allow image-conditioned sampling in latent space. This generative approach unlocks the exploration of mask completion or inpainting. The experimental validation on COCO and ADE20k yields strong segmentation results. Finally, we demonstrate our model’s adaptability to multi-tasking by introducing learnable task embeddings. The code and models will be made available. [^1] https://github.com/segments-ai/latent-diffusion-segmentation
Keywords:
Latent Diffusion Panoptic Segmentation Dense Prediction
@@ 1 Introduction
The image segmentation task [49, 54] has gained popularity in the literature, encompassing three popular subfields: semantic, instance, and panoptic segmentation. Over the years, segmentation tools have proven their usefulness for a wide range of applications, such as autonomous driving [19], medical imaging [52], agriculture [17], and augmented reality [1, 26]. Current methods are built upon convolutional networks [30] and transformers [77, 21] to learn hierarchical image representations, and to simultaneously leverage large-scale datasets [91, 47]. Earlier segmentation approaches relied on specialized architectures [29, 8, 72, 81], such as region proposal networks [63] and dynamic convolutions [35]. More recent approaches [15, 9] advocate for an end-to-end strategy but introduce complex loss functions, e.g., bipartite matching. Some works have shown promising results without the necessity for labels [74, 75, 76, 18, 80, 82], but likewise require highly-specialized modules, such as region proposal networks or clustering. Differently, we seek to leverage generative models to bypass the aforementioned components.
@@
Figure 1: (Left:) We present a simple generative approach for panoptic segmentation that builds upon Stable Diffusion [64]. The key idea is to leverage the diffusion process to bypass complex detection modules and to unlock mask inpainting. The generative process is conditioned on RGB images to iteratively predict the masks. (Right:) Our framework can be extended to a multi-task setting by introducing task embeddings.
This observation aligns with recent works [51, 10, 43, 11, 12] advocating for general computer vision models in an attempt to unify the field. These pioneering works limit the adoption of task-specific components but instead use a generative process. In a similar vein, we take inspiration from recent text-to-image diffusion models [64, 61, 56, 20, 32, 31, 62, 55] to tackle the segmentation task in a generative fashion. In addition to the diffusion model’s general architecture, we further motivate this decision as follows: (i) diffusion models generate images with high photorealism and diversity; (ii) perform on par with autoregressive priors while being more computationally efficient [68, 61]; (iii) learn high-quality spatial representations, advantageous for dense prediction tasks; (iv) naturally exhibit image-editing capabilities. We now aim to build upon these properties.
@@
To realize this objective, we introduce LDMSeg, a simple Latent Diffusion Model for Segmentation, visualized in Figure 1. Our contributions are fourfold:
@@
  • Generative Framework: We propose a fully generative approach based on Latent Diffusion Models (LDMs) for panoptic segmentation. We build upon Stable Diffusion [64] to strive for simplicity and computational efficiency.
@@
  • General-Purpose Design: As a result, we circumvent specialized architectures, complex loss functions, and object detection modules, present in the majority of prevailing methods. Here, the denoising objective omits the necessity for object queries, region proposals, and Hungarian matching [44] to handle the permutation-invariance of the instances.
@@
  • Mask Inpainting: We apply our method to scene-centric datasets. In contrast to prior art, we demonstrate mask inpainting for different sparsities.
  • Multi-Task Framework: Our simple and general approach can easily be extended to train a single generative model for multiple tasks, like instance segmentation, semantic segmentation and depth prediction. Querying the model for a different task merely requires changing the task embedding.
To the best of our knowledge, this paper presents the first latent diffusion approach that achieves strong results for panoptic segmentation, while also extending to mask inpainting and multi-task learning.
@@ 2 Related Work
Panoptic Segmentation. Panoptic segmentation [40] has lately gained popularity as it combines semantic and instance segmentation. In particular, its goal is to detect and segment both stuff-like (e.g., vegetation, sky, mountains, etc.) and thing-like (e.g., person, cat, car, etc.) categories. Earlier works modified instance segmentation architectures to additionally handle (amorphous) stuff categories as they are hard to capture with bounding boxes. For instance, Kirillov et al. showed promising results by making independent predictions using semantic and instance segmentation architectures [40], and later by integrating a semantic segmentation branch with a feature pyramid network (FPN) into Mask R-CNN [39]. Other works [85, 14] extend this idea by relying on specialized architectures and loss functions from both segmentation fields. More recent works [9, 78, 15, 16, 87, 90] handle things and stuff categories in a unified way via object queries and Hungarian matching [44]. These works generally depend on specialized modules – such as anchor boxes [63], non-maximum suppression [33, 7], merging heuristics [40, 85], or bipartite matching algorithms [9, 15, 78], etc. – to generate panoptic masks. Instead, we propose a task-agnostic generative framework to bypass these components. Consequently, we refrain from using task-specific augmentations, such as large-scale jittering or copy-paste augmentations [25].
@@
General-purpose Frameworks. Similar to our work, a few task-agnostic solutions have been suggested that cast vision problems as a generative process. Kolesnikov et al. [43] minimized task-specific knowledge by learning task-specific guiding codes with an LLM, to train a separate vision model. Chen et al. [12] simplified this procedure by framing vision tasks as language modeling tasks, within a single model. Lu et al. [51] also follow this route and show promising results for a large variety of vision and language tasks using a unified framework. Each work presents the input as a sequence of discrete tokens which are subsequently reconstructed via autoregressive modeling. Other works [79, 4, 55] leverage masked image modeling to train a single model for multiple vision tasks. Differently, we leverage the denoising process in continuous latent space, which is well-suited to handle dense prediction tasks with high-dimensional inputs [31, 10, 64].
@@
Denoising Diffusion Models. Denoising diffusion models [67, 31, 68] were introduced as a new class of generative models. Recent strategies [20, 56, 32, 61] additionally leverage text as guidance, e.g., via CLIP embeddings, to achieve results with impressive realism and control. Building upon its success, a few diffusion-based solutions appeared in the segmentation literature. However, they have undesirable properties: (i) the inability to differentiate between instances [2, 3, 5, 34, 83], (ii) the necessity for specialized architectures and loss functions [27, 86], or (iii) the dependence on object detection weights and bit diffusion [10, 13]. The closest related work from Chen et al. [10], presents a framework for panoptic segmentation by leveraging the diffusion process in pixel-space. In contrast, we rigorously follow latent diffusion models by relying on continuous latent codes without the necessity for object detection data. Importantly, we can effortlessly leverage public image-diffusion weights [64] as we keep the architecture task-agnostic.
@@ 3 Method
Figure 2: Overview of LDMSeg. Inspired by latent diffusion models, we present a simple diffusion framework for segmentation and mask inpainting. The approach consists of two stages: (i) learn continuous codes $z_{t}$ with a shallow autoencoder on the labels (Sec. 3.1); (ii) learn a denoising function conditioned on image latents $z_{i}$ (Sec. 3.2). In the second stage, the error between the predicted noise $\hat{\epsilon}$ and the applied Gaussian noise $\epsilon$ is minimized. During inference, we traverse the denoising process by starting from Gaussian noise. The models $f_{t}$ and $f_{i}$ respectively encode the labels and images. While we rely on the image encoder $f_{i}$ from Stable Diffusion [64], we focus on $f_{t}$ and $g$ for segmentation. We aim to prioritize generality by limiting task-specific components.
Preliminaries. This paper aims to train a fully generative model for panoptic segmentation via the denoising diffusion paradigm [31]. While this generative approach results in longer sampling times than discriminative methods, we justify this decision through four different lenses: (i) ease of use – diffusion models omit specialized modules, such as non-maximum suppression or region proposal networks, are stable to train, and exhibit faster sampling than autoregressive models [73]; (ii) compositionality – this generative approach captures complex scene compositions with high realism and diversity, while also enabling image editing [61]; (iii) dataset-agnostic – we rely on spatially structured representations that are not tied to predefined classes or taxonomies; (iv) computational cost – latent diffusion models [64] reduce the computational requirements by modeling latents instead of pixels with an autoencoder [38, 57]. Motivated by these points, we create a framework that builds upon latent diffusion models’ generative power [64, 61] for the segmentation task.
Problem Setup. The considered segmentation task requires a dataset of images $\mathcal{X}=\{x_{1},\ldots,x_{n}\}$ and corresponding ground truth panoptic segmentation masks $\mathcal{Y}=\{y_{1},\ldots,y_{n}\}$. We don’t make any distinction between things or stuff classes but handle all classes identically. We start by focusing on predicting panoptic IDs without the classes, rendering the method class-agnostic. Later, we extend our approach to also include class labels via task embeddings.
@@
From a high-level perspective, our method has two key components: learning the prior over segmentation latents $z_{t}$ and learning a conditional diffusion process in latent space. First, we train a shallow autoencoder to capture the prior distribution $p(z_{t})$ that learns to compress the labels into compact latent codes $z_{t}$. Second, we train a diffusion process – conditioned on the image and target features $p(y|z_{t},z_{i})$. This component is responsible for decoding noisy target features, guided by image features $z_{i}$. We make a similar derivation as [61] but condition the generative process on images. Formally, this two-step procedure allows us to construct the conditional distribution $p(y|x)$ via the chain rule as $p(y|x)=p(y|z_{t},z_{i})\cdot p(z_{t})$. Both terms, $p(z_{t})$ and $p(y|z_{t},z_{i})$, are reflected in our framework as two separate training stages (see Figure 2). In particular, Section 3.1 discusses the learning of the prior via autoencoding and Section 3.2 leverages this prior to train a latent diffusion model.
3.1 Stage 1: Compress Targets
In the first step, we train a network to compress the task-specific targets into latent codes. While we will focus on panoptic segmentation masks, we note that a similar strategy is applicable to other dense prediction targets (see Section 4.2).
Motivation. The motivation to design a shallow autoencoder stems from the observation that segmentation maps differ fundamentally from images as they are lower in entropy. First, these masks typically contain only a small number of unique values as they only capture the object’s general shape and location in the scene. Second, neighboring pixels are strongly correlated and often identical, resulting in largely spatially redundant information. We conclude that the segmentation task only necessitates a shallow network to efficiently compress the task’s targets and to reliably capture the prior distribution $p(z_{t})$. We hypothesize that this observation holds for a myriad of dense prediction tasks, such as panoptic segmentation, depth prediction, saliency estimation etc. This also justifies why we refrain from using more advanced autoencoders that rely on computationally demanding architectures with adversarial or perceptual losses [36, 89], e.g., VQGAN [23], typically used to encode images [64].
Encoding. We analyze several encoding strategies to represent the input segmentation map $y$, latents $z_{t}$ and output $\hat{y}$.
1. Input $y$: Let $N$ denote the maximum number of instances per image. RGB-encoding (3 channels), bit-encoding ($\log_{2}N$ channels), one-hot-encoding ($N$ channels), or positional-encoding [77] seem all justifiable options. However, one-hot encoding unnecessarily expands the input dimensions and bit encoding avoids the need for a fixed color palette for semantic segmentation. Hence, we follow [10] in representing the instances as bits.
@@ 3.1 Stage 1: Compress Targets
Architectural Design. The architecture of the autoencoder follows a simple and shallow design. It comprises an encoder $f_{t}$ and decoder $g$, yielding the overall function $g\circ f_{t}$. The encoder $f_{t}$ includes only a few strided convolutions to compress the targets and is inspired by ControlNet [88]. For instance, targets with size 512$\times$512 can be resized efficiently to 64$\times$64, in order to leverage the latent space of Stable Diffusion [64] by stacking 3 convolutions of stride 2. Similarly, the decoder $g$ consists of one or more transpose convolutions to upscale the masks and minimize a loss at pixel-level. As a consequence, the number of trainable parameters is at least 2 orders of magnitudes smaller than the amount of parameters in the diffusion model $h$ ($\approx 2$M vs. $800$M). The model’s shallow design brings several advantages to the table: fast training, good generalization across datasets, and applicability to inpainting without architectural changes.
Loss Function. The autoencoder aims to minimize the reconstruction error $\mathcal{L}_{\text{rec}}$ between the outputs $\hat{y}$ and the one-hot encoded segmentation masks $y$. While regression losses are a valid choice, we opt for a categorical loss due to the segmentation task’s discrete nature. The reconstruction loss comprises two loss terms: (i) the cross-entropy loss $\mathcal{L}_{\text{ce}}$ enforces unique and confident predictions for each pixel; (ii) the mask loss $\mathcal{L}_{\text{m}}$ further refines the segmentation masks by treating each instance individually, alleviating the need for exhaustively labeled images, in contrast to prior generalists [79, 43]. This term is implemented via the BCE and Dice losses [71, 29]. Note that this autoencoding strategy prevents the need for Hungarian matching [44].
Typically, latent diffusion models incorporate a penalty term $\Omega$ into the loss formulation to align the bottleneck latents with a standard Gaussian distribution $\mathcal{N}(0,I)$. This term generally takes the form of a KL divergence, resulting in a variational autoencoder [38]. However, we empirically observed that weight decay regularization suffices to keep the weights $w$, and by extension the latents $z_{t}$, bounded (see Section 4.3). Consequently, the magnitude of the model’s weights $||w||_{2}$ is penalized, resulting in the final loss formulation: